An envelope is 4cm longer than it is wide. The area is 96cm (with exponent of 2). Fine the length and width.
Answer by ikleyn(52781)

.

Let w be the width; then the length is (w+4).


The area equation is

    w*(w+4) = 96.


The standard form is

    w^2 + 4w - 96 = 0.


Factor left side

    (w+12)*(w-8) = 0


The roots are  -12 and 8;  but only positive value is the solution to the problem.


ANSWER.  The width is 8 inches,  the length is 8+4 = 12 inches.


CHECK.   8*12 = 96.    ! Correct !

Solved.

-------------

If you want to see many other similar problems, look into the lesson
    - Problems on the area and the dimensions of a rectangle
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic
"Dimensions and the area of rectangles and circles and their elements".

Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.